OpenStudy (anonymous):

10x^2-29x+10, find the factors. Is it prime? I can't think of any factors that would work in this situation.

hero (hero):

It is not prime

OpenStudy (anonymous):

what are the factors? i can't think of any.

hero (hero):

25*4

hero (hero):

10x^2 - 29x + 10 = 0 10x^2 - 25x - 4x + 10 = 0 5x(2x - 5) -2(2x-5) = 0 (2x-5)(5x-2) = 0

OpenStudy (anonymous):

thank you, SOOOOOOOOOOOOOOOOO much. You're a life saver.

hero (hero):

You still have no clue how to find the factors